### v2

@@ 4.3 Ablation Studies
We have done several ablation studies to show the contributions and properties of our proposed HAIR.
Effectiveness of Classifier and Hyper Selecting Net (HSN) As detailed in Table 4, we study the impact of the proposed modules in four settings: (a) Here, the model aligns with the Restormer architecture [12]. (b) In this configuration, the Global Information Vector (GIV) is directly used as conditional embeddings for the Transformer Blocks in the Decoder, rather than for weight generation. (c) The GIVs are designated as independently trainable parameters, each randomly initialized, with the Decoder’s Transformer Blocks having their own distinct GIVs. (d) This setup incorporates both components. The results demonstrate the indispensability of both components. With the addition of only 3M parameters and no change to the logical structure, Res-HAIR outperforms Restormer by 1.7 dB in PSNR, underscoring the simplicity and effectiveness of our proposed approach.
Position of Classifier (and HyperTrans Blocks) In our design approach, the Classifier is positioned after the first three Encoder levels, with the subsequent four Decoder levels employing HyperTrans Blocks, constituting a 3+4 configuration. The rationale for this design choice is demonstrated in Table 5. It is evident that the optimal placement for the Classifier is at the network’s midpoint. This positioning takes advantage of the expansive receptive field of the features post the network’s halfway point, while allowing the HSN to dynamically generate parameters for the remaining modules. Furthermore, it is crucial to strike a balance between the receptive field size of the features fed into the Classifier and the count of HyperTrans Blocks utilized, ensuring the model’s efficiency and adaptability.
Table 4: Impact of key components. Results are from single deraining task on Rain100L.
Setting Prams. Classifier HSN PSNR SSIM
(a) (baseline) 26M - - 36.74 .978
(b) 27M - 36.88 .979
(c) 28M - 36.76 .979
@@
Table 5: Impact of position. Results are from Denoise task on Urban100 ($\sigma$=25).
@@
HAIR for Other Baselines. We have previously posited that HAIR is essentially a plug-in-and-play module which is readily integrable with any image-to-image network architecture. To substantiate this claim, we have implemented HAIR on various established baselines. As depicted in Table Table 6, we selected three efficacious image restoration models—Transweather [1], AirNet [20], and Restormer [12]—for integration with HAIR. Specifically, we have integrated the Classifier at the network’s midpoint for each method and transitioned the subsequent layers to Hypernetworks-based modules. The outcomes show that our method outperforms PromptIR as a plug-in-and-play module, underscoring the universality and potency of our approach.
Table 6: HAIR for different baselines architectures. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images. Best performances are highlighted.
[…]
Transweather+PromptIR 51M 22.89 .920 29.79 .913 29.95 .877 25.74 .781 23.02 .849 26.28 .868
Transweather+HAIR 42M 23.66 .935 32.34 .947 29.96 .875 26.33 .802 23.16 .858 27.09 .884
AirNet+PromptIR 14M 21.34 .883 33.52 .953 30.92 .882 24.37 .786 18.18 .737 25.67 .848
AirNet+HAIR 10M 22.15 .899 34.56 .957 30.94 .884 25.44 .792 18.24 .740 26.27 .854
Restormer+PromptIR 36M 30.61 .974 36.17 .973 31.25 .887 27.93 .851 22.89 .842 29.77 .905
Restormer+HAIR 29M 30.62 .978 38.11 .981 31.49 .891 28.52 .874 23.12 .847 30.37 .914
@@ 5 Conclusion
This paper introduces HAIR, a novel plug-and-play Hypernetworks-based module capable of being easily integrated and adaptively generating parameters for different networks based on the input image. Our method comprises two main components: the Classifier and the Hyper Selecting Net (HSN). Specifically, the Classifier is a simple image classification network with Global Average Pooling, designed to produce a Global Information Vector (GIV) that encapsulates the global information from the input image. The HSN functions as a Fully-connected Neural Network, receiving the GIV and outputting parameters for the corresponding modules. Extensive experiments indicate that HAIR can significantly enhance the performance of various image restoration architectures at a low cost without necessitating any changes to their logical structures. By incorporating HAIR into the widely recognized Restormer architecture, we have achieved State-Of-The-Art performance on a range of image restoration tasks. The potential for further exploitation of data-conditioned Hypernetworks in image-to-image tasks is substantial, considering their robust adaptability to diverse input images based on their content.

### v4

@@ 4.3 Ablation Studies
We have done several ablation studies to evaluate the effectiveness of our proposed HAIR.
Effectiveness of DAC and HSN. As detailed in Tab. 5, we study the impact of the proposed modules in four settings: (a) The model aligns with the Restormer architecture [60]. (b) In this configuration, the Global Information Vector (GIV) is directly used as conditional embeddings for the Transformer Blocks in the Decoder, rather than for weight generation. (c) The GIVs are designated as independently trainable parameters, each randomly initialized, with the Decoder’s Transformer Blocks having their distinct GIVs. (d) This setup incorporates both components. The results demonstrate the indispensability of both components. With the addition of only 3M parameters and no change to the logical structure, Res-HAIR outperforms Restormer by 1.7 dB in PSNR, demonstrating its simplicity and effectiveness.
Table 4: HAIR for different baseline architectures. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images. Best performances are highlighted.
[…]
Transweather [48]+PromptIR [38] 51M 22.89 .920 29.79 .913 29.95 .877 25.74 .781 23.02 .849 26.28 .868
Transweather [48]+HAIR 42M 23.66 .935 32.34 .947 29.96 .875 26.33 .802 23.16 .858 27.09 .884
AirNet [27]+PromptIR [38] 14M 21.34 .883 33.52 .953 30.92 .882 24.37 .786 18.18 .737 25.67 .848
AirNet [27]+HAIR 10M 22.15 .899 34.56 .957 30.94 .884 25.44 .792 18.24 .740 26.27 .854
Restormer [60]+PromptIR [38] 36M 30.61 .974 36.17 .973 31.25 .887 27.93 .851 22.89 .842 29.77 .905
Restormer [60]+HAIR 29M 30.62 .978 38.11 .981 31.49 .891 28.52 .874 23.12 .847 30.37 .914
Position of DAC and HyperTrans Blocks. In our approach, the DAC is positioned after the first three Encoder levels, with the subsequent four Decoder levels employing HyperTrans Blocks, constituting a 3+4 configuration. The rationale for such design is shown in Tab. 6. As can be seen, the optimal placement for the DAC is at the network’s midpoint. This positioning takes advantage of the expansive receptive field of the features post the network’s halfway point, while allowing the HSN to dynamically generate parameters for the remaining modules. Furthermore, it is crucial to strike a balance between the receptive field size of the features fed into the DAC and the count of HyperTrans Blocks utilized, ensuring the model’s efficiency and adaptability.
Table 5: Impact of the existence of DAC and HSN. Results are from single deraining task on Rain100L.
Setting Prams. DAC HSN PSNR SSIM
(a) (baseline) 26M 36.74 .978
(b) (with DAC) 27M 36.88 .979
(c) (with HSN) 28M 36.76 .979
@@
Table 6: Impact of position of DAC and HyperTrans Blocks. Results are from Denoise task on Urban100 ($\sigma$=25).
@@
HAIR for Different Baselines. We have previously posited that HAIR is essentially a plug-in-and-play module which is readily integrable with any existing network architecture. To substantiate this claim, we have implemented HAIR on various baselines. As depicted in Tab. 4, we selected three efficacious image restoration models, i.e. Transweather [48], AirNet [27], and Restormer [60] for integration with HAIR. Specifically, we have integrated the DAC at the network’s midpoint for each method and transitioned the subsequent layers to Hypernetworks-based modules. The results show that our HAIR can significantly improve the performance of these baselines. Additionally, our HAIR outperforms PromptIR as a plug-in-and-play module, demonstrating its better application value.
@@ 5 Conclusion
This paper introduces HAIR, a novel plug-and-play Hypernetworks-based module capable of being easily integrated and adaptively generating parameters for different networks based on the input image. Our method comprises two main components: the Degradation-Aware Classifier (DAC) and the Hyper Selecting Net (HSN). Specifically, the DAC is a simple image classification network with Global Average Pooling, designed to produce a Global Information Vector (GIV) that encapsulates the global information from the input image. The HSN functions as a fully-connected neural network, receiving the GIV and outputting parameters for the corresponding modules. Extensive experiments indicate that HAIR can significantly enhance the performance of various image restoration architectures at a low cost without necessitating any changes to their logical structures. By incorporating HAIR into the widely recognized Restormer architecture, we have achieved SOTA performance on a range of image restoration tasks.
Supplementary Material
6 Detailed Proof of Theorem 1, Theorem 2
Notations We consider $\mathcal{X}=[-1,1]^{m_{1}}$ and $\mathcal{I}=[-1,1]^{m_{2}}$ and denote, $m:=m_{1}+m_{2}$. Here $\mathcal{X}$ stands for the set of input images $\mathbf{X}$, meanwhile $\mathcal{I}$ refers to the set of degradation information $e(\mathbf{X})$. For a closed set $X\subset\mathbb{R}^{n}$, we denote by $C^{r}(X)$ the linear space of all $r$-continuously differentiable functions $h:X\to\mathbb{R}$ on $X$ equipped with the supremum norm $\|h\|_{\infty}:=\max_{x\in X}\|h(x)\|_{1}$.
Given that vector functions can be complex, we define a scalar function $f^{i}:\mathbb{R}^{M}\xrightarrow{}\mathbb{R}(i=1,\cdots M)$ of a vector function $f:\mathbb{R}^{M}\xrightarrow{}\mathbb{R}^{M}$. Since $f^{i}$ is the $i$-th element of $f$, then we have
$=\min_{g}\max_{i}\max_{\mathbf{X}}|g^{i}(\mathbf{X})-y^{i}(\mathbf{X})|$
$=\min_{g}\max_{\mathbf{X}}|g^{t(g)}(\mathbf{X})-y^{t(g)}(\mathbf{X})|.$
Since the value range of $i$ is limited, given the function $g$, we can always find an integer $t(g)$ to replace $\max_{i}$. To complete the proof, we need some assumptions.
Assumption 1.
The target function $y^{i}\in\mathcal{W}_{r,3HW},i=1,2,\cdots 3HW$. The Sobolev space $\mathcal{W}_{r,3HW}$ is the set of functions that are $r$-times differentiable ($r\geq 1$) with all $r$-th order derivatives being continuous and bounded by the Sobolev norm $\leq 1$, defined on $\mathbb{R}^{3HW}$.
Intuitively, since $y^{i}$ maps the degraded image to a single pixel of the clean image, a slight disturbance in the degraded image should only bring a slight difference to the single pixel, so $y^{i}$ can be smooth and differentiable, or at least continue. Moreover, the domain and range of $y$ are restricted to $[0,1]^{3HW}$, with most pixels far less than 1, the Sobolev norm of $y_{i}$ can generally be less than 1 and thus this assumption generally holds.
Assumption 2.
For various functions (e.g. network) $g$, the integer $t(g)$ in Eq. 9 remains the same.
In our context, $g$ represents the same network with different weights during training. No matter how the parameters update, the most "difficult" degraded image generally remains the same one, e.g., the one with very severe degradation. Within this "difficult" image, the most "difficult" pixel should also remain the same, e.g. the pixel that varies dramatically from its clean version. In this way, we simplify the vector function $g$ to scalar function $g^{t(g)}$. With the two assumptions and all the assumptions in [15], we can directly use Theorem 3,4 and 5 to analyze the complexity.
Theorem 3.
(The Theorem 2 in [15]) Let $\sigma$ be a universal, piece-wise $C^{1}(\mathbb{R})$ activation function with $\sigma^{\prime}\in BV(\mathbb{R})$ and $\sigma(0)=0$. Let $\mathcal{E}_{\mathbf{e},\mathbf{q}}$ be a neural embedding method. Assume that $\mathbf{e}$ is a class of continuously differentiable neural network $e$ with zero biases, output dimension $k=\mathcal{O}(1)$ and $\mathcal{C}(e)\leq\ell_{1}$ and $\mathbf{q}$ is a class of neural networks $q$ with $\sigma$ activations and $\mathcal{C}(q)\leq\ell_{2}$. Let $\mathbb{Y}:=\mathcal{W}_{1,m}$. Assume that any non-constant $y\in\mathbb{Y}$ cannot be represented as a neural network with $\sigma$ activations. If the embedding method achieves error $d(\mathcal{E}_{\mathbf{e},\mathbf{q}},\mathbb{Y})\leq\epsilon$, then, the complexity of $\mathbf{q}$ is: $N_{\mathbf{q}}=\Omega\left(\epsilon^{-(m_{1}+m_{2})}\right)$.
The notation $BV(\mathbb{R})$ stands for the set of functions of bounded variation,
$BV(\mathbb{R}):=\left\{f\in L^{1}(\mathbb{R})\mid\|f\|_{BV}<\infty\right\}\textnormal{ where, }$, (10)
$\|f\|_{BV}:=\sup\limits_{\begin{subarray}{c}\phi\in C^{1}_{c}(\mathbb{R})\\ \|\phi\|_{\infty}\leq 1\end{subarray}}\int_{\mathbb{R}}f(x)\cdot\phi(x)\;\textnormal{d}x$
Note that a distinct neural network $e$ is not mandatory. For example, the "prompts" in PromptIR [38] are a set of trainable parameters that do not require a separate network to generate them. Yet, the conclusion remains the same even if network $e$ is non-existent.
Theorem 4.
(The Theorem 3 in [15]) In the setting of Theorem 3, except $k$ is not necessarily $\mathcal{O}(1)$. Assume that the first layer of any $q\in\mathbf{q}$ is bounded $\|W^{1}\|_{1}\leq c$, for some constant $c>0$. If the embedding method achieves error $d(\mathcal{E}_{\mathbf{e},\mathbf{q}},\mathbb{Y})\leq\epsilon$, then, the complexity of $\mathbf{q}$ is: $N_{\mathbf{q}}=\Omega\left(\epsilon^{-\min(m,2m_{1})}\right)$.
Theorem 5.
(The Theorem 4 in [15]) Let $\sigma$ be as in Theorem 3. Let $y\in\mathbb{Y}=\mathcal{W}_{r,m}$ be a function, such that, $y_{I}$ cannot be represented as a neural network with $\sigma$ activations for all $I\in\mathcal{I}$. Then, there is a class, $\mathbf{g}$, of neural networks with $\sigma$ activations and a network $f(I;\theta_{f})$ with ReLU activations, such that, $h(x,I)=g(x;f(I;\theta_{f}))$ achieves error $\leq\epsilon$ in approximating $y$ and $N_{\mathbf{g}}=\mathcal{O}\left(\epsilon^{-m_{1}/r}\right)$.
In the realm of image restoration, $m_{1}$ equals $3HW$, and $m_{2}$ equals $k$, where $k$ denotes the dimensionality of the flattened degradation embedding. In our method, $k$ is consistent with the shape of the Global Information Vector (GIV), specifically $2C$, and thus is $\mathcal{O}(1)$. Conversely, in PromptIR [38], $k$ is dynamic and contingent on the resolution of the input image, precluding it from being $\mathcal{O}(1)$. Theorem 5 indicates that the complexity for a Hypernetworks-based method to attain an error of $\epsilon$ is $\mathcal{O}\left(\epsilon^{-3HW/r}\right)$. Theorems 3 and 4 collectively suggest that the complexity for embedding-based methods is at least $\Omega\left(\epsilon^{-\min(3HW+k,6HW)}\right)$. This comparison shows that Hypernetworks-based methods like HAIR may require fewer parameters to reach a given error threshold compared to their embedding-based counterparts.
7 Discussion
7.1 HAIR for Conflicting Degradations.
As previously discussed in the Introduction (Sec. 1), the performance of conventional All-in-One image restoration methods, which rely on a single model with static parameters, can be significantly compromised when dealing with conflicting image degradations. To demonstrate this, we trained models on various combinations of datasets, each representing different types of degradations, and subsequently evaluated these models on their respective benchmarks. The outcomes are presented in Tables 7 and 8.
From a frequency domain analysis perspective, haze is identified as low-frequency noise, whereas rain and Gaussian noise are categorized as high-frequency disturbances. Conflicts arise when the degradations in a combined scenario include both low- and high-frequency noise components. According to Table 7, it is evident that the presence of conflicting degradations, such as the combination of noise and haze or rain and haze, can severely degrade the model’s performance. In contrast, when the combined degradations do not conflict, like the combination of noise and rain, the performance loss is minimal and may even enhance the overall performance.
Table 8 exhibits a similar trend, but with a notably reduced impact from conflicting degradations. This reduction in performance impairment underscores the effectiveness of our proposed Hypernetworks-based approach, which dynamically generates parameters based on the input image’s content. This adaptability allows our method to mitigate the performance loss typically associated with conflicting degradations.
Table 7: Performance of the PromptIR [38], when trained on different combinations of degradation types (tasks) i.e., removal of gaussian noise, rain and haze. Note that the "combination" here stands for combination of Datasets with single degradations instead of composite degradations. "Conflicting" here shows if the combined degradations are conflicting to each other. PSNR/SSIM are reported.
Degradation Conflicting Denoising on BSD68 dataset Deraining on Dehazing on
Noise Rain Haze $\sigma=15$ $\sigma=25$ $\sigma=50$ Rain100 SOTS
- 34.34/0.938 31.71/0.898 28.49/0.813 - -
- - - - 37.04/0.979 -
- - - - - 31.31/0.973
34.26/0.937 31.61/0.895 28.37/0.810 39.32/0.986 -
33.69/0.928 31.03/0.880 27.$7$ 30.09/0.975
- - - 35.12/0.969 30.36/0.973
33.98/0.933 31.31/0.888 28.06/0.799 36.37/0.972 30.58/0.974
Table 8: Performance of the our proposed Res-HAIR, when trained on different combinations of degradation types (tasks) i.e., removal of gaussian noise, rain and haze. PSNR/SSIM are reported.
Degradation Conflicting Denoising on BSD68 dataset Deraining on Dehazing on
Noise Rain Haze $\sigma=15$ $\sigma=25$ $\sigma=50$ Rain100 SOTS
- 34.36/0.938 31.72/0.898 28.50/0.813 - -
- - - - 39.00/0.985 -
- - - - - 31.68/0.980
34.33/0.937 31.66/0.896 28.44/0.811 41.55/0.989 -
34.13/0.935 31.49/0.892 28.22/0.803 - 31.18/0.979
- - - 38.44/0.983 31.22/0.979
34.16/0.935 31.51/0.892 28.24/0.803 38.59/0.983 30.98/0.979
7.2 HAIR for Unseen Composite Degradation.
Since HAIR can accurately discriminate unseen composite degradations, as illustrated in Fig. 2(d), we test Res-HAIR with these settings, as shown in Table 9, 11, 10. The results somehow shows the generalization ability of our proposed method. However, we consider the restoration outcomes for such degradations are not satisfactory enough. The interesting fact is, sometimes HAIR generates GIVs that are intermediate when confronted with composite degradations. This tendency results in weights that are a midpoint between those associated with each individual degradation type. For example, an image with both noise and haze degradations results in weights that are intermediate between the weights for images affected by noise alone and those affected by haze alone, which fails to fully eliminate either degradation and result in a somehow unsatisfactory performance. This insight reveals HAIR’s operational mechanism, highlighting its strategy for weights generation.
Table 9: Results of unseen composite degradation (haze + noise) on SOTS.
Method PSNR SSIM
PromptIR 16.211 0.785
Res-HAIR 16.798 0.802
Table 10: Results of unseen composite degradation (rain + noise) on Rain100L.
Method PSNR SSIM
PromptIR 24.348 0.726
Res-HAIR 24.365 0.729
Table 11: Results of unseen composite degradation (rain + haze) on SOTS.
Method PSNR SSIM
PromptIR 23.784 0.686
Res-HAIR 23.823 0.692
8 Possible Confusions
8.1 Does HAIR Also Utilize Fixed Parameters as Previous Methods Do?
In the Introduction (Sec. 1), we highlighted that our proposed HAIR method differs from previous approaches, which typically employ a static set of parameters to handle various image degradations. In contrast, HAIR employs a data-conditioned Hypernetwork to dynamically generate parameters based on the input image’s content. However, a potential point of confusion arises from the fact that the Hypernetworks in HAIR (i.e., the Hyper Selecting Nets) are indeed fixed during inference, suggesting that HAIR also relies on a set of fixed parameters to address different degradations.
To clarify this confusion, it is crucial to understand our motivation for using Hypernetworks. Our goal is to mitigate the performance loss caused by conflicting degradations. For example, by generating distinct parameters, we aim to enable the model to function either as a low-frequency or high-frequency filter, depending on the input image’s requirements. As illustrated in Fig. 1, the primary focus should be on the main networks that directly process the image.
Therefore, when we refer to "fixed parameters" and "dynamic parameters," we are referring to the parameters of the main network that interacts with the image, not the Hypernetworks responsible for parameter generation, which do not directly engage with the input image. The point lies in the dynamic adaptation of the main network’s parameters to the specific characteristics of the input image, which is the innovative aspect of HAIR.
9 More Implementation Details
Figure 5: Overview of the Transformer block used in the Res-HAIR framework. The Transformer block is composed of two sub modules,the Multi Dconv head transposed attention module(MDTA) and the Gated Dconv feed-forward network(GDFN).
9.1 Transformer Block in Res-HAIR Framework
(This section is directly borrowed from the paper of PromptIR. [38]) In this section, we present the block diagram5 of the transformer block and further, elaborate on the details of the transformer block employed in the Res-HAIR framework. The transformer block follows the design and hyper-parameters outlined in [60]
To begin, the input features $\mathbf{X}\in\mathbb{R}^{H_{l}\times W_{l}\times C_{l}}$ are passed through the MDTA module. In this module, the features are initially normalized using Layer normalization. Subsequently, a combination of $1\times 1$ convolutions followed by $3\times 3$ depth-wise convolutions are applied to project the features into Query ($\mathbf{Q}$), Key ($\mathbf{K}$), and Value ($\mathbf{V}$) tensors. An essential characteristic of the MDTA module is its computation of attention across the channel dimensions, rather than the spatial dimensions. This effectively reduces the computational overhead. To achieve this channel-wise attention calculation, the $Q$ and $K$ projections are reshaped from $H_{l}\times W_{l}\times C_{l}$ to $H_{l}W_{l}\times C_{l}$ and $C_{l}\times H_{l}W_{l}$ respectively, before calculating dot-product, hence the resulting transposed attention map with the shape of $C_{l}\times C_{l}$. Bias-free convolutions are utilized within this submodule. Furthermore, attention is computed in a multi-head manner in parallel.
After MDTA Module the features are processed through the GDFN module. In the GDFN module, the input features are expanded by a factor $\gamma$ using $1\times 1$ convolution and they are then passed through $3\times 3$ convolutions. These operations are performed through two parallel paths and the output of one of the paths is activated using GeLU non-linearity. This activated feature map is then combined with the output of the other path using element-wise product.
9.2 Implementation Details of HyperTrans Blocks
This section elaborates on specific details presented in Sec. 3.1.2. Within the Restormer framework [60], each Transformer consists of a total of 12 logical convolutional layers, as depicted in Fig. 5. For code implementation, we utilize 6 convolutional layers to effectively emulate the functionality of 12 convolutional layers within each Transformer Block. Consequently, the generated parameters $\mathbf{w}\in\mathbb{R}^{P}$ described in Sec. 3.1.2 encompass all the parameters necessary for these 6 convolutional layers of a single Transformer Block. The Weights Box $\mathbf{W}\in\mathbb{R}^{N\times P}$, therefore, represents the aggregate parameters for $N$ Transformer Blocks.
Given a fixed Global Information Vector (GIV) $\mathbf{V_{g}}$, it is processed through the Fully-Connected Neural Network (FCNN) followed by a Softmax operation to yield the Selecting Vector $\mathbf{V_{s}}\in\mathbb{R}^{N}$. This vector is subsequently employed to "select" the definitive parameters from the Weights Box. Each Decoder level possesses a distinct Weights Box, which is shared among all the HyperTrans Blocks at that level. It is crucial to highlight that all HyperTrans Blocks across all levels operate using the same GIV derived from a single Classifier. Meanwhile, each single HyperTrans Block is equipped with its own Hyper Selecting Net, resulting in a unique Selecting Vector and, consequently, its own set of parameters for each.
10 Visual Results
In Fig. 6 and Fig. 7 we provide more visual results to show the strong ability of our method for image restoration tasks.
Figure 6: Additional visual results.
Figure 7: Additional visual results